# FutureVision: A methodology for the investigation of future cognition

### v1

@@ Conclusion
Our central finding that the FutureVision methodology allows for assessing the impact of base space violations on cognitive load is an important contribution to several fields of reasearch where future cognition plays a central role. Beyond its contributions to Scenario Forecasting, Speculative Futures and Quantitative Futurism, the methodology presented in this paper can inform research in other fields such as human-computer interaction, cognitive neuroscience, and clinical decision-making. Understanding how individuals process counterfactual and valence-based scenarios can support the development of adaptive systems that respond to users’ cognitive states, such as virtual reality environments for cognitive training or rehabilitation.
Data and code availability
Data and scripts are available via https://osf.io/9kvu2

### v2

@@ Conclusion
Our central finding that the FutureVision methodology allows for assessing the impact of base space violations on cognitive load is an important contribution to several fields of research where future cognition plays a central role. Beyond its contributions to Scenario Forecasting, Speculative Futures and Quantitative Futurism, the methodology presented in this paper can inform research in other fields such as human-computer interaction, cognitive neuroscience, and clinical decision-making. Understanding how individuals process counterfactual and valence-based scenarios can support the development of adaptive systems that respond to users’ cognitive states, such as virtual reality environments for cognitive training or rehabilitation.